40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1783     Entry time: Thu Jul 23 10:05:38 2009
Author: Alberto 
Type: Update 
Category: PSL 
Subject: Summary of the latest adventures with the alignment of the mode cleaner 

Alberto, Koji,

Summary of the latest adventures with the alignment of the mode cleaner

Prior events.

  • Last week, on July 12th the RFM network crashed (elog entry 1736). I don't know for sure which one was the cause and which one the effect, but also C1DAQADW was down and it didn't want to restart. Alex fixed it the day after.
  • On the evening of Sunday July 20th I noticed that the mode cleaner was unlocked. A closer inspection showed me that MCL was frozen at -32768 counts. To fix that I rebooted C1DCUEPICS and burtrestored to snapshots from the day before.
  • On Tuesday July 21st another failure of the RFM Network made necessary a reboot of the frame builder and of all front end computers (entry 1772). As a consequence, the mode cleaner couldn't get locked anymore, even if the mirror's sliders in the MC-Align MEDM screen were in the proper positions. At that time I missed to check the MC suspension positions as a way to ensure that the MC hadn't really changed. Although later, as it turned out, that would have been useless anyway since all the data record prior to the computers crash of that day somehow had been corrupted (entry 1774). Neither the MC2 LSC control or MC ASC control could engage so I (erroneously) thought that some tune of the periscope might help. So I did but, since the Mode Cleaner was misaligned, that had the effect of spoiling the good matching of the periscope to the MC cavity.
  • Yesterday, Wednesday July 22nd, I found out about the sticky slider effect (entry 1776). At that point we didn't have anymore a way to know that the MC optics were actually in their proper original alignment state because of the lack of a reference for those in the data record (as I wrote above). I had to go back to the periscope and fix the alignment.


Chronicles of periscope and MC alignment

Yesterday morning I started aligning the periscope but it turned out to be trickier than usual. With the ASC (Alignment Sensing Control) off and only the length controls on, the Mode Cleaner didn't lock easily, although I knew I wasn't very far from the sweet spot.

In the afternoon the struggle continued and the matching of the the beam to the MC cavity became just worse. At some point I noticed that the ASC inputs somehow had got on - although the ASC still looked disabled from the MClock MEDM main screen. So I was actually working against the Wave Front Sensors and further worsening the periscope alignment.

That hurled me to the weeds. After hours of rowing across the stormy waters of a four-dimensional universe I got to have occasional TEM00 flashes at the transmission but still, surprisingly, no MC locking. Confused, I kept tuning the periscope but that just kicked me off road again.

Then at about 7pm Koji came to my rescue and suggested a more clever and systematic way to solve the problems. He suggested to keep record of the MC mirrors alignment state and re-align the cavity to the periscope. Then we would gradually bring the cavity back to the original good position changing the periscope alignment
at the same time.

 

That would have worked straight away, if we hadn't been fighting against a subtle and cruel enemy: the 40m computer network. But I (as John Connor), and Koji (as the Terminator) didn't pull back.

Here's a short list of the kinds of weapons that the computers threw to us:

  1. After a while the FSS entered a funny state. It showed transmission: we had light at the MC (and even flashes) but the MEDM readout of the FSS transmitted power after the cavity was low (~0.019). Also the spot on the monitor showed a slightly different pattern from how I remembered it. On the other side the transmission camera didn't show that typical halo as usual.
  2. MCL was frozen at 32768. I ran the MCDown and MCUp script a couple of times and that unstuck it.
  3. On op340m we found that the MC autolocker script wasn't running. So I restarted it. Still nothing changed: bright and sharp flashes appeared on the monitor (sign of a not too bad alignment) but no lock.
  4. I rebooted C1IOO. No change.
  5. I rebooted C1DCUEPICS and burtrestored the EPICS computers to Jul 19th. No change.
  6. Then I burtrestored the c1psl.snapshot and that finally did something. The FSS reflected spot changed and the halo appeared again at its transmission camera. Soon after the MC got locked.


We then proceeded with Koji's plan. In an iterative process, we aligned the MC cavity maximizing the transmission and tuned the periscope in order to match the Faraday input of the interferometer. The last thing we did it by looking at the camera pointing at the Faraday isolator.

We found that we didn't have to tune the periscope much. That means that all afternoon I didn't really go too far, but the autolocker wasn't working properly, or it wasn't working at all.

Then we ran the alignment script for the X arm but it didn't work before we aligned the steering mirrors.

Then we ran it three times but could not get more than 0.87 at TRX. That means that there we still have to work on the alignment to the Faraday. That's job for today in the trenches of the lab.

 

ELOG V3.1.3-